OpenStudy (anonymous):

Consider the function f(x) =(x^2-x-12)/(x-4) . Describe the graph of this function. Include all discontinuities, intercepts, and the basic shape of the graph.

OpenStudy (anonymous):

\[\frac{x^2-x-12}{x-4}=x+3 \]A graph of f(x) = x + 3 is a line. A plot is attached.

OpenStudy (anonymous):

Thank you (:

OpenStudy (anonymous):

You're welcome.